OpenStudy (anonymous):

find the product of this expression (a - 12)(a + 12) =

OpenStudy (anonymous):

Have you learned how to multiply two binomials? hint: FOIL Method

OpenStudy (cggurumanjunath):

(a+b)(a-b)=a^2-b^2 a=a b=12